OpenStudy (anonymous):

two triangles left = perimeter =14 right=perimeter=21 Y Y 5/4Y 5/4Y X 3X (sorry, i cant put it the triangles)

OpenStudy (anonymous):

okay so you can solve this by creating equations for triangle 1: lets say that perimeter (p) = 2y + x =14

OpenStudy (anonymous):

and triangle 2: p = 2 (5/4 y) = x

OpenStudy (anonymous):

we can manipulate the equation for triangle 1 to be x= 14 -2y. then substitute that into the equation for triangle 2 so the equation would be: 5/2y + 3( 14-2y) =21

OpenStudy (anonymous):

k i got the first part but now i dont get how you got 5/2y+3(14-2y)=21

OpenStudy (mathstudent55):

Left: x + y + y = 14 Right: 3x + (5/4)y + (5/4)y = 21 x + 2y = 14 3x + (10/4)y = 21 or 3x + (5/2)y = 21 Multiply 2nd eq by 2: 6x + 5y = 42 Now rewrite the two equations: x + 2y = 14 6x + 5y = 42 multiply top eq by -6, and write second eq below: -6x - 12y = -84 6x + 5y = 42 Now add the equations: -7y = -42 y = 6 Now substitute y = 6 into original first eq: x + 2y = 14 x + 2(6) = 14 x + 12 = 14 x = 2 Answer: x = 2, y = 6

OpenStudy (anonymous):

Thank you!